### v1

# MA-LMM: Memory-Augmented Large Multimodal Model for Long-Term Video Understanding

Abstract
With the success of large language models (LLMs), integrating the vision model into LLMs to build vision-language foundation models has gained much more interest recently. However, existing LLM-based large multimodal models (e.g., Video-LLaMA, VideoChat) can only take in a limited number of frames for short video understanding. In this study, we mainly focus on designing an efficient and effective model for long-term video understanding. Instead of trying to process more frames simultaneously like most existing work, we propose to process videos in an online manner and store past video information in a memory bank. This allows our model to reference historical video content for long-term analysis without exceeding LLMs’ context length constraints or GPU memory limits. Our memory bank can be seamlessly integrated into current multimodal LLMs in an off-the-shelf manner. We conduct extensive experiments on various video understanding tasks, such as long-video understanding, video question answering, and video captioning, and our model can achieve state-of-the-art performances across multiple datasets.
1 Introduction
Large language models (LLMs) have gained significant popularity in the natural language processing field. By pre-training on large-scaled textual data, LLMs (e.g. GPT [1, 2, 3, 4], LLaMA [5, 6]) have demonstrated remarkable abilities to perform both generative and discriminative tasks with a unified framework. Recently, there has been a growing interest in utilizing LLMs on multimodal tasks. By integrating LLMs with visual encoders, they can take images and videos as input and show incredible capabilities in various visual understanding tasks, such as captioning, question answering [7, 8, 9, 10, 11, 12, 13], classification, detection, and segmentation [14, 15, 16, 17, 18, 19, 20].
Figure 1: (a) We propose the long-term memory bank to auto-regressively store and accumulate past video information, different from previous methods directly feeding the visual encoder’s outputs into the querying transformer. (b) GPU memory consumption v.s. video frame length of existing multimodal methods and MA-LMM during inference. Circle sizes represent the number of text tokens.
To handle video inputs, some prior large multimodal models [7, 9] directly feed the concatenated query embeddings of each frame along the temporal axis into LLMs. However, the inherent context length limitation of LLMs and GPU memory consumption restrict the number of video frames that can be processed. For example, LLaMA has a context length limitation of 2048 while large multimodal models like LLaVA [8] and BLIP-2 [7, 9] take in 256 and 32 tokens per image respectively. Therefore, this design is not practical and feasible when video duration is much longer (e.g. movies and TV shows). To address these issues, a naive solution is to apply average pooling along the temporal axis like Video-ChatGPT [21], but this leads to inferior performances as it lacks explicit temporal modeling. An alternative method involves adding a video modeling component to capture temporal dynamics, as seen in Video-LLaMA [12], which employs an extra video querying transformer (Q-Former) to obtain video-level representation. However, this design adds model complexities, increases the training parameters, and is not suitable for online video analysis.
With these in mind, we introduce a Memory-Augmented Large Multimodal Model (MA-LMM), aiming for efficient and effective long-term video modeling. MA-LMM adopts a structure similar to existing large multimodal models [7, 9, 12], which comprise a visual encoder to extract visual features, a querying transformer to align the visual and text embedding spaces, and a large language model. As illustrated in Figure 1(a), as opposed to directly feeding visual encoder outputs to the querying transformer, we opt for an online processing approach that takes video frames sequentially and stores the video features in the proposed long-term memory bank. This strategy of sequentially processing video frames and leveraging a memory bank significantly reduces the GPU memory footprint for long video sequences. It also effectively addresses the constraints posed by the limited context length in LLMs as demonstrated in Figure 1(b). Our design provides a solution for long-term video understanding with large multimodal models with great advantages over prior approaches [7, 9, 12, 21, 13] which consume huge GPU memory and require a large number of input text tokens.
The core contribution of our approach is the introduction of a long-term memory bank that captures and aggregates historical video information. Specifically, the memory bank aggregates past video features in an auto-regressive manner, which can be referenced during subsequent video sequence processing. Also, our memory bank is designed to be compatible with the Q-Former, where it acts as the key and value in the attention operation for long-term temporal modeling. As a result, it can be seamlessly integrated into existing large multimodal models in an off-the-shelf manner to enable long-term video modeling ability. To further enhance efficiency, we propose a memory bank compression method that maintains the length of the memory bank constant relative to the input video length. By selecting and averaging the most similar adjacent frame features, it can preserve all the temporal information while significantly reducing the temporal redundancies in long videos.
We summarize our main contributions as follows:
  • We introduce a novel long-term memory bank design to enhance existing large multimodal models, equipping them with long-term video modeling capability.
  • Our model significantly reduces the GPU memory usage and addresses LLMs’ context length limitations by processing video sequences in an online fashion.
  • Our approach has achieved new state-of-the-art performances on various downstreaming video tasks, including long-term video understanding, video question answering, and video captioning.
2 Related Work
Image-language models. Inspired by the success of powerful language models [4, 5, 6, 1, 2, 3], recent image-language models tend to incorporate pre-trained language models with image encoders to support the multimodal reasoning ability [22, 7, 8, 9, 10]. Flamingo [22] proposes to connect powerful pre-trained vision-only and language-only models and achieve state-of-the-art performance in few-shot learning tasks. BLIP-2 [7] introduces a lightweight querying transformer to bridge the modality gap between the frozen pre-trained image encoder and frozen LLMs. Despite having significantly fewer trainable parameters, it performs well on various multimodal tasks. LLaVA [8] employs a simple linear layer to project image features into the text embedding space and efficiently finetunes LLMs  [23] for better performance. Building upon BLIP-2, MiniGPT-4 [10] collects a large-scale high-quality dataset of image-text pairs and achieves better language generation ability. VisionLLM [15] leverages the reasoning and parsing capacities of LLMs, producing strong performance on multiple fine-grained object-level and coarse-grained reasoning tasks.
Video-language models. Current image-language models such as Flamingo [22] and BLIP-2 [7, 9] can also support video inputs. They simply flattened the spatio-temporal features into 1D sequences and then fed them into the language models for video inputs. However, these approaches can not effectively capture the temporal dynamics of videos. Recent video-language models [11, 12, 13] attempt to address this by incorporating mechanisms to better understand these dynamics. For instance, Video-LLaMA [12] enhances BLIP-2 structure by adding an additional video querying transformer to explicitly model the temporal relationship. Similarly, building on LLaVA [8], Video-ChatGPT [21] simply average pools the frame-level features across spatial and temporal dimensions to generate video-level representation. ChatVideo [24] treats tracklets as the basic video elements and relies on several video foundation models to annotate these tracklets with textual descriptions. Then, it uses ChatGPT [4] to achieve various video understanding tasks. VideoChat [13] utilizes perception models to generate action and object annotations, which are then forwarded to LLMs for further reasoning. Despite these advancements, these models are primarily designed for short videos, as longer videos pose significant challenges due to language model context length and GPU memory restrictions.
Figure 2: (a) Framework overview. MA-LMM auto-regressively processes video frames in an online manner. Two long-term memory banks are designed to store the raw visual features and learned queries at each timestep, which are used for future reference. The Q-Former is composed of several cascaded blocks, indexed by $l$. LLM outputs text for various video understanding downstream tasks. The snowflake icon indicates components with fixed parameters, while the flame icon denotes parts of the model that are fine-tuned. (b) Illustration of the memory bank compression technique, which is applied to maintain the length of the memory bank constant.
Long-term video models. Long-term video understanding methods focus on capturing long-range patterns in long videos, which typically exceed 30 seconds. To mitigate the computational demands of processing long videos, a prevalent approach involves using pre-extracted features, sidestepping the need for joint training of backbone architectures [25, 26, 27, 28, 29]. Alternatively, some research works aim to devise sparse video sampling methods [30, 31], reducing the number of input frames by only preserving salient video content. Other works like Vis4mer [32] and S5 [33] leverage the streamlined transformer decoder structure of S4 [34] to enable long-range temporal modeling with linear computation complexity. Inspired by the memory bank design [35, 36, 37, 38], we propose to integrate the long-term memory bank with large multimodal models to enable efficient and effective long-term temporal modeling capabilities.
3 Method
We introduce MA-LMM, a memory-augmented large multimodal model for long-term video understanding. Figure 2(a) illustrates the overview of our MA-LMM framework. Following similar practices of large multimodal models [7, 9, 8, 12], the overall model architecture can be divided into three parts: (1) visual feature extraction with a frozen visual encoder (Sec. 3.1), (2) long-term temporal modeling with a trainable querying transformer (Q-Former) to align the visual and text embedding spaces (Sec. 3.2), and (3) text decoding with a frozen large language model (Sec. 3.3).
3.1 Visual Feature Extraction
We propose to auto-regressively process video frames in an online manner. This design draws inspiration from the cognitive processes humans use to handle long-term visual information. Instead of concurrently processing extensive duration of signals, humans process them in a sequential manner, correlate current visual inputs with past memories for comprehension, and selectively retain salient information for subsequent reference [38]. Similarly, our MA-LMM processes video frames sequentially, dynamically associating new frame input with historical data stored in the long-term memory bank, ensuring that only discriminative information is conserved for later use. This selective retention facilitates a more sustainable and efficient approach to video understanding, which further allows the model to automatically support online video reasoning tasks.
Formally, given a sequence of $T$ video frames, we pass each video frame into a pre-trained visual encoder and obtain the visual features $V=[v_{1},v_{2},..,v_{T}],v_{t}\in\mathbb{R}^{P\times C}$, where $P$ is the number of patches for each frame and $C$ is the channel dimension for the extracted frame feature. Then we inject temporal ordering information into the frame-level features by a position embedding layer ($PE$) as
$f_{t}=v_{t}+PE(t),f_{t}\in\mathbb{R}^{P\times C}.$ (1)
3.2 Long-term Temporal Modeling
For aligning the visual embedding to the text embedding space, we use the same architecture as the Querying Transformer (Q-Former) in BLIP-2 [7, 9]. Q-Former takes in the learned queries $z\in\mathbb{R}^{N\times C}$ to capture video temporal information, where $N$ is the number of learned queries, and $C$ is the channel dimension. In our experiments, Q-Former outputs 32 tokens for each image, which is more efficient than 256 tokens produced by LLaVA [8]. Each Q-Former block consists of two attention submodules: (1) cross-attention layer, which interacts with the raw visual embedding extracted from the frozen visual encoder, and (2) self-attention layer, which models interactions within the input queries. Different from the original Q-Former in BLIP-2 that only attends to the current frame’s embedding, we design a long-term memory bank consisting of the visual memory bank and the query memory bank, which accumulates the past video information and augments the input to cross- and self-attention layers for effective long-term video understanding.
Visual Memory Bank. The visual memory bank stores the raw visual features of each frame extracted from the frozen visual encoder. Specifically, for the current time step $t$, the visual memory bank contains the concatenated list of past visual features $F_{t}=\texttt{Concat}[f_{1},f_{2},..,f_{t}],F_{t}\in\mathbb{R}^{tP\times C}$. Given the input query $z_{t}$, the visual memory bank acts as the key and value as:
$Q$ $=z_{t}W_{Q},\ K=F_{t}W_{K},\ V=F_{t}W_{V}.$ (2)
Then we apply the cross-attention operation as:
$O$ $=Attn(Q,K,V)=\text{Softmax}\left(\dfrac{QK^{T}}{\sqrt{C}}\right)V.$ (3)
In this way, it can explicitly attend to past visual information through the cached visual memory bank with long-term context. Since all the cross-attention layers in the Q-Former attend to the same visual feature, there is only one visual memory bank that is shared across all the Q-Former blocks.
Query Memory Bank. Different from the fixed visual memory bank which stores the raw and static visual features, the query memory bank accumulates input queries of each timestep, represented as $Z_{t}=\texttt{Concat}[z_{1},z_{2},..,z_{t}],Z_{t}\in\mathbb{R}^{tN\times C}$. By storing these queries, we maintain a dynamic memory of the model’s understanding and processing of each frame up to the current timestep via the Q-Former. The query memory bank also acts as key and value as:
$Q$ $=z_{t}W_{Q},\ K=Z_{t}W_{K},\ V=Z_{t}W_{V}.$ (4)
similar to the Eq 2. Then we apply the same attention operation as Eq. 3. At each time step, $z_{t}$ contains the learned important information specifically for each video till the current timestep $t$. Different from the static visual memory bank, the input queries $z_{t}$ evolve through cascaded Q-Former blocks during the model training, capturing distinct video concepts and patterns at increasing levels of abstraction. As a result, each self-attention layer has a unique query memory bank, where the contained input queries are updated during the training time.
Memory Bank Compression. Given that our model directly stores past video information in the memory banks, the GPU memory cost and computational complexity increase linearly as the number of past video frames. This becomes particularly challenging for long videos, and thus it is essential to further compress the memory bank to a smaller size. A straightforward solution is to use a first-in-first-out queue to manage the temporal sequence. In this scenario, the features of the earliest timestep are removed if the current length of the memory bank exceeds a pre-defined limit. However, this approach inevitably leads to the loss of earlier historical information as new frames are added and old features are popped when the memory bank is full. To address this issue, we take advantage of the temporal redundancy in long videos and introduce a novel memory bank compression (MBC) technique. This method aggregates and compresses video information over time based on the similarity of adjacent features, while preserving early historical information. In this manner, repetitive information is consolidated in the memory bank, while discriminative features remain intact.
We apply the compression algorithm at each auto-regressive iteration if the current memory bank length is larger than the pre-defined threshold $M$. Formally, given the visual memory bank containing a list of $f_{1},f_{2},..,f_{M}],f_{t}\in\mathbb{R}^{P\times C}$, when a new frame feature $f_{M+1}$ comes in, we need to compress the memory bank by reducing the length by 1. At each spatial location $i$, we first calculate the cosine similarity between all the temporally adjacent tokens as
$s^{i}_{t}=\cos(f^{i}_{t},f^{i}_{t+1}),t\in[1,M],i\in[1,P].$ (5)
Then we select the highest similarity across time, which can be interpreted as the most temporally redundant features:
$k=\mathrm{argmax}_{t}(s^{i}_{t}).$ (6)
Next, we simply average the selected token features at all the spatial locations to reduce the memory bank length by 1:
$\hat{f}^{i}_{k}=(f^{i}_{k}+f^{i}_{k+1})/2.$ (7)
In this way, we can still preserve the most discriminative features while keeping the temporal ordering unchanged as depicted in Figure 2(b). The same procedure is adopted to compress the query memory bank.
3.3 Text Decoding
As we process video frames in an auto-regressive manner, the Q-Former output at the final timestep contains all historical information, which is then fed into the LLM. Therefore, we can significantly reduce the number of input text tokens from NTN*T to $N$, addressing the context length limitation of the current LLMs and substantially easing the GPU memory requirements. During training, given a labeled dataset consisting of video and text pairs, our model is supervised with the standard cross entropy loss as:
$\mathcal{L}=-\frac{1}{S}\sum_{i=1}^{S}\log P(w_{i}|w_{<i},V).$ (8)
in which $V$ represents the input video, and $w_{i}$ is the $i$-th ground-truth text token. During training, we update the parameters of the Q-Former while keeping the weights of both the visual encoder and the language model frozen.
4 Experiments
Table 1: Comparison with state-of-the-art methods on the LVU [29] dataset. Bold and underline represent the top-1 and top-2 results.
Model Content Metadata Avg
Relation Speak Scene Director Genre Writer Year
Obj_T4mer [29] 54.8 33.2 52.9 47.7 52.7 36.3 37.8 45.0
Performer [39] 50.0 38.8 60.5 58.9 49.5 48.2 41.3 49.6
Orthoformer [40] 50.0 38.3 66.3 55.1 55.8 47.0 43.4 50.8
VideoBERT [41] 52.8 37.9 54.9 47.3 51.9 38.5 36.1 45.6
LST [32] 52.5 37.3 62.8 56.1 52.7 42.3 39.2 49.0
VIS4mer [32] 57.1 40.8 67.4 62.6 54.7 48.8 44.8 53.7
S5 [33] 67.1 42.1 73.5 67.3 65.4 51.3 48.0 59.2
Ours 58.2 44.8 80.3 74.6 61.0 70.4 51.9 63.0
Table 2: Comparison on the Breakfast [42] and COIN [43] datasets. The top-1 accuracy is reported here.
Model Breakfast COIN
TSN [44] - 73.4
VideoGraph [45] 69.5 -
Timeception [28] 71.3 -
GHRM [46] 75.5 -
D-Sprv. [47] 89.9 90.0
ViS4mer [32] 88.2 88.4
S5 [33] 90.7 90.8
Ours 93.0 93.2
4.1 Tasks and Datasets
To validate the effectiveness of the proposed MA-LMM, we mainly focus on the long-term video understanding task. We also extend the evaluation to standard video understanding tasks (e.g., video question answering, video captioning) to further compare with existing multimodal methods.
Long-term Video Understanding. We conduct experiments on three widely used long-term video datasets including LVU [29], Breakfast [42], and COIN [43]. We report the top-1 classification accuracy as the evaluation metric. The LVU dataset contains $\sim$30K videos extracted from $\sim$3K movies, with each video lasting 1 to 3 minutes. Given that current large multimodal models generally perform text generation and lack regression capability, we limit our experiments to seven classification tasks: relationship, speaking style, scene, director, genre, writer, and release year. The Breakfast [42] dataset includes videos related to breakfast preparation, which consists of 1712 videos with an average length of around 2.7 minutes. COIN [43] is a large-scale dataset for comprehensive instructional video analysis, which comprises 11827 instructional videos from YouTube, covering 180 distinct tasks in 12 domains related to daily life. The average length of a video is 2.36 minutes.
Video Question Answering. We conduct evaluation on three open-ended video question answering datasets including MSRVTT-QA [48], MSVD-QA [48], and ActivityNet-QA [49]. ActivityNet-QA contains long videos with average durations of 2 minutes, while MSRVTT-QA and MSVD-QA consist of short videos with 10-15 seconds duration.
Video Captioning. We report the video captioning results of METEOR [50] and CIDEr [51] metrics on three popular datasets: MSRVTT [52], MSVD [53] and Youcook2 [54].
Online Action Prediction. We further evaluate the online prediction capability of our model by conducting experiments on the EpicKitchens-100 [55] dataset, which consists of 700 long videos of cooking activities with 100 total hours. It includes 97 verbs, 300 nouns, and 3807 action types. Following the same experimental setting in  [56], we report the top-5 accuracy and recall results on the validation dataset.
4.2 Implementation Details
For the visual encoder, we adopt the pre-trained image encoder ViT-G/14 [57] from EVA-CLIP [58], it can be further changed to other clip-based video encoders. We use the pre-trained Q-Former weights from InstructBLIP [9] and adopt Vicuna-7B [59] as the LLM. All the experiments are conducted on 4 A100 GPUs. More details about training and evaluation are described in the supplementary material.
Table 3: Comparison with state-of-the-art methods on the video question answering task. Top-1 accuracy is reported.
Model MSRVTT MSVD ActivityNet
JustAsk [60] 41.8 47.5 38.9
FrozenBiLM [61] 47.0 54.8 43.2
SINGULARITY [62] 43.5 44.1
VIOLETv2 [63] 44.5 54.7
GiT [64] 43.2 56.8
mPLUG-2 [65] 48.0 58.1
UMT-L [66] 47.1 55.2 47.9
VideoCoCa [67] 46.3 56.9 56.1
Video-LLaMA [12] 46.5 58.3 45.5
Ours 48.5 60.6 49.8
Table 4: Comparison with state-of-the-art methods on the video captioning task. METEOR (M) and CIDEr (C) results are reported.
Model MSRVTT MSVD YouCook2
M C M C M C
UniVL [68] 28.2 49.9 29.3 52.8 127.0
SwinBERT [69] 29.9 53.8 41.3 120.6 15.6 109.0
GIT [64] 32.9 73.9 51.1 180.2 17.3 129.8
mPLUG-2 [65] 34.9 80.3 48.4 165.8
VideoCoca [67] 73.2 128.0
Video-LLaMA 32.9 71.6 49.8 175.3 16.5 123.7
Ours 33.4 74.6 51.0 179.1 17.6 131.2
4.3 Main Results
Long-term Video Understanding. We compare MA-LMM with previous state-of-the-art (SOTA) methods on the LVU benchmark [29] in Table 2. Notably, MA-LMM outperforms existing long-term video models (S5 [33], ViS4mer [32], VideoBERT [41], and Object Transformer [29]) in both content understanding and metadata prediction tasks. This results in significant improvement in most tasks, enhancing the average top-1 accuracy by 3.8% compared to the S5 [33] model. Unlike previous video-based models which process all video frames simultaneously in an offline manner and predict probabilities for each class, our MA-LMM processes video frames in an online fashion and directly outputs the text label for each class type.
We also evaluate our MA-LMM on the Breakfast [42] and COIN [43] datasets that pose a challenge for the long-term video activity classification task. We show the results in Table 2. Our method improves upon the previous best method, S5[33], by 2.3% and 2.4% respectively on the top-1 accuracy metric. This result further proves the superior long-term video understanding capability of our approach.
Video Question Answering. To compare with existing multimodal video understanding methods, we conduct experiments on the open-ended video question answering datasets in Table 4 to demonstrate the generalization ability of our model. Given that these are mostly short videos, it is expected that our memory bank will be less effective. Interestingly, we observe that our MA-LMM achieves new state-of-the-art performances on the MSRVTT and MSVD datasets while falling short of VideoCoCa’s performance on the ActivityNet dataset. On the latter, it is not surprising, since VideoCoCa [67] leverages large-scale video-text datasets for pre-training (e.g., HowTo100M [70] and VideoCC3M [71]) while our MA-LMM uses model weights only pre-trained on the image-text datasets.
Notably, our MA-LMM significantly outperforms the recent LLM-based model Video-LLaMA [12] on all three datasets. Video-LLaMA concatenates all the query embeddings from the frozen image Q-Former and trains an additional video Q-Former from scratch to model temporal dependencies, consuming too much GPU memory to be feasible for long video inputs. In contrast, our MA-LMM simply fine-tunes the weights from the pre-trained image Q-Former without introducing an additional video Q-Former, yet is able to effectively capture temporal relationships by virtue of the long-term memory bank. This result strongly justifies the superiority of our design on the general video question answering task, and reveals that even a few frames and queries captured in the memory banks can have significant beneficial effects.
Video Captioning. To further evaluate the capabilities of our MA-LMM in generating free-form text, we conduct experiments on the standard video captioning datasets including MSRVTT [52], MSVD [53] and YouCook2 [54] in Table 4. Although these datasets only consist of videos with short duration and our model is initially pre-trained merely on image-text dataset pairs, our MA-LMM exhibits outstanding performances across all the metrics. It consistently ranks among the top-2 positions compared to current leading methods. Remarkably, our results also surpass the recent Video-LLaMA [12] on these datasets, highlighting the significant improvements our model offers in both video captioning and question-answering tasks.
Online Action Prediction. Since our model can naturally support the online video understanding task, we compare our MA-LMM with Video-LLaMA on the EpicKitchens-100 [55] dataset to investigate the online action prediction capability. In Table 5, our MA-LMM outperforms Video-LLaMA, achieving more accurate results in both top-5 accuracy and recall measures. This highlights our model’s superior capacity to anticipate actions in an online manner, showcasing its effectiveness for applications that require real-time analytical capabilities.
Table 5: Action anticipation results on EpicKitchens-100.
Model Accuracy@Top-5 Recall@Top-5
Verb Noun Act. Verb Noun Act.
Video-LLaMA 73.9 47.5 29.7 26.3 27.3 11.7
Ours 74.5 50.7 32.7 25.9 29.9 12.2
Table 6: Contribution of visual and query memory banks.
Visual Query LVU Breakfast COIN
48.3 74.6 72.3
61.5 91.8 92.4
58.0 81.4 88.5
63.0 93.0 93.2
Table 7: Contribution of the long-term memory bank (MB) under off-the-shelf evaluation without training.
MB MSRVTT MSVD ActivityNet LVU
19.5 38.8 29.9 23.6
20.3 40.0 37.2 32.8
Table 8: Ablation of different temporal modeling methods.
Method #Frame #Token GPU LVU Breakfast COIN
Concat 60 1920 49.2 62.6 90.4 93.0
Avg Pool 100 32 21.2 57.6 80.6 87.6
ToMe 100 200 22.2 61.5 91.3 91.5
FIFO 100 32 19.1 61.3 88.5 90.4
MBC 100 32 19.1 63.0 93.0 93.2
Table 9: The comparison of using different LLMs.
LLM MSRVTT MSVD ActivityNet LVU
FlanT5-XL 46.5 57.6 48.2 62.0
Vicuna-7B 48.5 60.6 49.8 63.0
4.4 Ablation Studies
Contribution of each component. To further investigate the contribution of the visual memory bank and query memory bank, we conduct ablation studies in Table 9. Initially, we observe that without any memory bank module, the performances across all three datasets are notably worse, due to the lack of temporal context. The introduction of either memory bank results in substantial improvements, confirming their roles in enhancing the model’s ability to understand temporal sequences. We also find that the visual memory bank achieves better performance than the query memory bank. We hypothesize that the explicit method of storing historical raw video features in the visual memory bank is more effective than the query memory bank which implicitly captures video information through the input learned queries. And two memory banks are complementary to each other. When incorporating two memory banks together, our approach can boost the final performance by 14.7%, 18.4%, and 20.9% on the LVU, Breakfast, and COIN, respectively.
Long-term temporal modeling ablation. We compare different temporal modeling approaches in Table 9. In our setup, the Q-Former outputs 32 text tokens per frame. The most straightforward approach for temporal feature integration is either concatenating or averaging frame-level features. However, they resulted in inferior performances. Notably, concatenation requires a significantly higher number of text tokens and computational cost compared to other variants, which also introduces higher GPU memory consumption since they need to takes in all the video frames simultaneously. In addition, we conduct experiments using ToMe [72] to reduce the number of text tokens per frame from 32 to 2. However, without our auto-regressive strategy, it still requires 200 text tokens for 100-frame input. The second part of this table presents the performances of different memory bank compression approaches. The first-in-first-out (FIFO) technique removes the oldest features to main the length of the memory bank fixed, while the memory bank compression (MBC) strategy merges temporally consecutive features with the highest similarity, effectively reducing the most redundant information while keeping the temporal ordering unchanged. With this design that theoretically keeps all historical information, MBC outperforms FIFO by 1.7%, 4.5%, and 2.8% accuracy across three datasets. This experimental result validates the superior efficiency and effectiveness of our approach in modeling long-term temporal information.
Off-the-shelf evaluation. A key advantage of MA-LMM is that our long-term memory bank can be inserted into existing large multimodal models in an off-the-shelf manner, thereby endowing them with effective temporal modeling capabilities without retraining. As presented in Table 9, MA-LMM can consistently boost the final performance when incorporating the long-term memory bank to the baseline method [9]. Particularly, on long-term video datasets like ActivityNet and LVU, MA-LMM can largely improve the results by 7.3% and 9.2%. This highlights the robustness of long-term memory banks in temporal modeling under the off-the-shelf setting.
Figure 3: Impact of different memory bank lengths.
Different language model architectures. Our MA-LMM can utilize different language model architectures including but not limited to encoder-decoder models and decoder-only models. We experimented with two popular models FlanT5-XL [73] and Vicuna-7B [59], and show the results in Table 9 that the Vicuna-7B marginally outperforms the FlanT5-XL on these video tasks.
Figure 4: Visualization results on the video question answering task and the online off-the-shelf setting.
Figure 5: Visualization of the compressed visual memory bank.
Memory bank length ablation. In Figure 3, we conduct experiments to evaluate the effect of varying the memory bank length. Given an input of 100 video frames, the top-1 accuracy first increases as the feature bank length becomes larger. This rise can be attributed to the augmented storage capacity of the memory bank, which can preserve more historical data and consequently boost the final performance. However, we observe that performances begin to saturate when the memory bank length is around 10 to 20. This supports our hypothesis that there are prevalent temporal redundancies in long videos, and we can significantly reduce the frame length without sacrificing the performance.
4.5 Visualization
In Figure 4, we provide a comprehensive visual comparison between MA-LMM and Video-LLaMA [12]. In the video question answering task, our MA-LMM exhibits superior memorization and recognition capabilities. Specifically, it can accurately memorize historical information and recognize fine-grained information, such as the color of the man with No.7, and precisely count the number of goalkeepers who appeared in the video. With the auto-regressive design, our model supports online reasoning directly. This capability is further exemplified in our experiments on off-the-shelf evaluations using custom questions. We can see that, our model can correctly anticipate the next step of the video ("egg will be cooked") and predict the correct recipe of the video ("scrambled egg"). More visualization examples are shown in the supplementary material.
Figure 5 provides a visualization of the compressed visual memory bank. We set the memory bank length to 5 for this illustration. The compressed visual memory bank appears to group consecutive frames with similar visual content. For instance, in the presented video, the video frames are effectively grouped into five clusters, each capturing a distinct yet semantically consistent activity, which is similar to the effect of temporal segmentation.
5 Conclusion
In this paper, we introduce a long-term memory bank designed to augment current large multimodal models, equipping them with the capabilities to effectively and efficiently model long-term video sequences. Our approach processes video frames sequentially and stores historical data in the memory bank, addressing LLMs’ context length limitation and GPU memory constraints posed by the long video inputs. Our long-term memory bank is a plug-and-play module that can be easily integrated into existing large multimodal models in an off-the-shelf manner. Experiments on various tasks have demonstrated the superior advantages of our method. We believe our MA-LMM offers valuable insights for future research in the long-term video understanding area.
Acknowledgements. This project was partially funded by NSF CAREER Award (#2238769) to AS.
Appendix
We report additional ablation experiments in Sec. A. We also present more qualitative results on the video captioning task in Sec. B. And in Sec. C, we show more dataset-specific implementation details and hyper-parameters. Finally, we discuss some limitations and future works in Sec. D.
Appendix A Additional Experiments
Memory bank compression at different spatial levels. In Table 10, we show comparison results of compressing the memory bank at different spatial levels (frame-level vs. token-level) on the LVU [29], Breakfast [42] and COIN [43] datasets. For the frame-level compression, we calculate the cosine similarity between adjacent frame features and average the frame-level features with the highest similarity. For the token-level compression, the cosine similarity is calculated between tokens at the same spatial location across the entire temporal axis, given that each frame-level feature contains multiple tokens at different spatial locations. The results indicate that token-level compression consistently surpasses frame-level compression in performance. Particularly, on the Breakfast dataset, the token-level surpasses the frame-level by 6.5% in top-1 accuracy. This superiority can be attributed to the importance of recognizing the object type of breakfast in videos. And token-level compression can help preserve much more fine-grained spatial information and details.
Inference time of different input frames In Figure 6, the inference time of MA-LMM increases linearly with respect to the frame lengths, due to its auto-regressive design of processing video frames sequentially. In contrast, directly concatenating frame-level features takes much longer time and higher GPU memory consumption, since it needs to process all video frames simultaneously.
Appendix B More Qualitative Results
Our model’s enhanced capabilities in video captioning are further showcased through additional visualization results in Figure 7. Here, our MA-LMM significantly outperforms Video-LLaMA [12] in generating detailed and accurate sentence descriptions. For instance, in the first video, our model precisely describes the action as "remove the onion rings and place them on the paper towel," capturing the entire action steps, while Video-LLaMA’s description lacks this completeness, notably missing the crucial action of removing the onion rings. In the second video example, our model distinguishes itself by accurately identifying subtle details such as specific ingredients: chili powder, salt, and garlic powder, which Video-LLaMA overlooks. This highlights the enhanced capability of our MA-LMM in recognizing and describing fine-grained details.
Appendix C Experiment Details
We show the details of hyper-parameters in the following table for different tasks and datasets. For all the experiments, we use a cosine learning rate decay. Table 11 shows the hyper-parameters for the long-term video understanding task. For the LVU dataset, we follow the same practice in [32, 33], we sample 100 frames of 1 fps for each video clip. For the Breakfast [42] and COIN [43], we uniformly sample 100 frames from the whole video. Table 12 shows the hyper-parameters on the MSRVTT-QA [48], MSVD-QA [48], and ActivityNet-QA [49] datasets for the video question answering task while Table 13 presents the hyperparameters on the MSRVTT [52], MSVD [53], YouCook2 [54] datasets for the video captioning tasks.
Table 10: Memory bank compression at different spatial levels.
Spatial Level LVU Breakfast COIN
Frame-level 61.8 86.5 91.1
Token-level 63.0 93.0 93.2
Figure 6: Inference time vs. input frame length.
Figure 7: Visualization results on the video captioning task.
Appendix D Limitation and Future Work
Since our model takes in video frames in an online manner, leading to reduced GPU memory usage, but at the cost of increased video processing time. This trade-off becomes particularly noticeable with extremely long videos, where processing times can become significantly prolonged. To mitigate this issue, we suggest a hierarchical method to process extremely long-term video sequences. This strategy involves dividing extensive videos into smaller segments and then processing each segment sequentially in an auto-regressive fashion as we present in the main paper. Then we can employ additional video modeling techniques to model inter-segment relationships. This method aims to strike a balance between memory efficiency and processing speed, making it a practical solution for long-term video understanding.
For the future work, there are several potential aspects to further enhance the model’s capabilities. First, replacing the existing image-based visual encoder with a video or clip-based encoder can naturally enhance the model’s ability to capture short-term video dynamics. This provides a better representation of the video’s temporal dynamics. Second, the model’s overall performance in understanding videos can substantially benefit from the pre-training stage on large-scale video-text datasets. This approach is a common practice in existing research and has proven effective in enhancing generalization capabilities. Finally, the flexibility inherent in our model’s architecture allows for the incorporation of a more advanced LLM as the language decoder. This integration offers a clear opportunity for boosting the final performance, making our model more effective in interpreting and responding to complex video content.
Table 11: Hyperparameters of different datasets on the long-term video understanding task.
Dataset LVU Breakfast COIN
LLM Vicuna-7B
Epochs 20
Learning rate 1e-4
Batch size 64
AdamW $\beta$ (0.9, 0.999)
Weight decay 0.05
Image resolution 224
Beam size 5
Frame length 100
Memory bank length 20
Prompt “What is the {task} of the movie?” “What type of breakfast is shown in the video?” “What is the activity in the video?”
Table 12: Hyperparameters of different datasets on the video question answering task.
Dataset MSRVTT MSVD ActivityNet
LLM Vicuna-7B
Epochs 5
Learning rate 1e-4
Batch size 128
AdamW $\beta$ (0.9, 0.999)
Weight decay 0.05
Image resolution 224
Beam size 5
Frame length 20
Memory bank length 10
Prompt “Question: {} Short Answer:”
Table 13: Hyperparameters of different datasets on the video captioning task.
Dataset MSRVTT MSVD YouCook2
LLM Vicuna-7B
Epochs 10
Learning rate 1e-5 1e-5 1e-4
Batch size 128
AdamW $\beta$ (0.9, 0.999)
Weight decay 0.05
Beam size 5
Image resolution 224
Frame length 80
Memory bank length 40
Prompt “what does the video describe?”